OpenStudy (anonymous):

Devin spent a total of $64 dollars for shoes and a hat. The shoes cost $6 more than the hat. How much did the shoes cost? Let shoes be represented by x and hats be represented by y. A. $38 B. $26 C. $35 D. $29

OpenStudy (opcode):

Set up a system of equations: \[64 = x + y \\ x = 6 + y\] \[64 = (6 + y) + y \\ 64 = 2y + 6 \\ 64 - 6 = 2y \\ 58 = 2y \\ 58 \div 2 = y \\ 29 = y \] Now that you know \(y = 29\) sub that into the original equation and solve for \(x\).

OpenStudy (anonymous):

So its d.

OpenStudy (opcode):

No, we want to find how much the shoes cost. \(y\) is the hats. Remember that the shoes are \(6\) more than the hats so: \(29+6 = x\)

OpenStudy (anonymous):

C.

OpenStudy (opcode):

Correct! To fact check this we do: \(35 + 29 = 64\) Which is true, so you are correct.

OpenStudy (anonymous):

yes well thank you do you think you can help me again please ?

OpenStudy (opcode):

I can try, just open a new thread.